OpenStudy (anonymous):

What is the average common ratio between the successive height values of ball 1? Ball 2? Experimental errors may cause common ratios to have some variances within the data for one ball. Use the average common ratio. @dan815

OpenStudy (anonymous):

ball 1 = 2.9 ball 2 = 2.08

OpenStudy (anonymous):

@Disco619

OpenStudy (disco619):

Im only good at certain kind of Geometry... srry :/

OpenStudy (anonymous):

lol this is algebra 2 jst pls help any thing will do @Disco619

OpenStudy (disco619):

That's what Im saying, im no good at Maths except some Geometry. I got nthin :/

OpenStudy (anonymous):

tussi yaar meri situation ko samjo. do you know anybody that can help me? @Disco619

OpenStudy (disco619):

Dekhta hoon... @robtobey ???

OpenStudy (anonymous):

yeh toh offline hai

OpenStudy (disco619):

@RadEn ?? @shamil98 ??

OpenStudy (shamil98):

What is the average common ratio between the successive height values of ball 1? Ball 2 is this asking for the common ratio between ball 1 and 2, or the attempts of ball 1 , if so you should have more attempts listed..

OpenStudy (anonymous):

it asking for the common ratio @shamil98

OpenStudy (shamil98):

between what?

OpenStudy (shamil98):

the successive height values of ball 1? There should be more than one value listed then..

OpenStudy (anonymous):

ok for ball one there are 2.9 ft 1.5 ft 1 ft 0.6 ft 0.46 ball 2: 2.08 1.25 ft 0.5 ft 0.25 stop bouncing

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

The common ratio is the number you multiply to get the next number , now it says that you will have variances with the common ratio so you are supposed to find the average of the ratios and use that..

OpenStudy (shamil98):

What i'm a bit unsure is that the ratios between the terms are quite different..

OpenStudy (shamil98):

2.9/1.5 = 1.93 1/0.6 = 1.67 0.6/0.46 = 1.3

OpenStudy (anonymous):

yeah it says to use the average common ratio

OpenStudy (shamil98):

yeah just find the average of those ratios that i put forth. add them up and divide by the number of terms (which is 3)

OpenStudy (shamil98):

find the common ratios for ball B the same way and average them the same way then

OpenStudy (anonymous):

What is the height of each ball on the fifth bounce (i.e., Height 6)? Use the geometric sequence formula, an = a1rn – 1 and show your work. @shamil98

OpenStudy (shamil98):

you have your average common ratio, your first term, so just plug it into the formula.

OpenStudy (anonymous):

then whats rn? @shamil98

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

rn = r*n r = ratio n = the number of the term you want to find out